OpenStudy (anonymous):

Need help!!! Determine the domain of the function (f*g)(x) where f(x) = 1/x , g(x) = sqrt(2-5x)

OpenStudy (anonymous):

is this \[(f\cdot g)(x) \text{ or } (f\circ g)(x)\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so this is a composition of the functions. so we'll get \[(f \circ g)(x)=f(g(x))=\frac{ 1 }{ g(x) }=\frac{ 1 }{ \sqrt{2-5x} }\] right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what has to happen?

OpenStudy (anonymous):

remember... you're not allowed to divide by 0 and you can't have a negative number inside a square root.

hero (hero):

The restriction is \(x \ne 0\) and \(x \ngeq \dfrac{2}{5}\) Furthermore, I would rationalize the denominator.